OpenStudy (anonymous):

how do you find the tan of 7pi/6?

OpenStudy (aravindg):

hmm..7 pi/6 =pi+pi/6

OpenStudy (anonymous):

First convert to degrees, that's 180 + 30 = 210

OpenStudy (anonymous):

Next, find the tan 210. Do you know how to do that?

OpenStudy (anonymous):

i don't see where you got pi+pi over 6?

OpenStudy (anonymous):

or how you converted that to degrees? this isn't making sense. I need to be able to do this without a calculator..

OpenStudy (aravindg):

tan(pi+pi/6)=-tan theta=-tan pi/6=1/root3

OpenStudy (anonymous):

Well, 7pi/6 is the same as pi, or 6pi/6 plus another pi/6 to get 7pi/6.

OpenStudy (aravindg):

value=1/root 3

OpenStudy (anonymous):

like i understand tan = cos/sin

OpenStudy (anonymous):